OpenStudy (anonymous):

What are the intercepts of the line x - 3y = 9?

OpenStudy (yanasidlinskiy):

set it in y=mx+b form:) 3y=x-9

OpenStudy (yanasidlinskiy):

If i'm not mistaking....

OpenStudy (anonymous):

y=(9,0) x=(0,-3)

OpenStudy (acxbox22):

x intercept is when y is 0 so, x-3(0)=9 x=9 <<<--- x int is (9,0)

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

Thanks @YanaSidlinskiy , @katyperryeioqe , and @acxbox22 ! :)))

OpenStudy (anonymous):

x=9 and y=-3 sorry

OpenStudy (acxbox22):

y intercept is when x is 0 so, (0)-3y=9 -3y=9 y=-3 <<<<---- y int is (0,-3)

OpenStudy (yanasidlinskiy):

No problem!!!!:) Glad I could help you!!!!!!!!!!!:)